prettygirl15:

help plz The expression 4(1/2)^x represents the height of a bouncing ball after x bounces. What does the value 4 represent in the expression?

Vocaloid:

Notice what happens when you plug in x = 0 4(1/2)^0 = 4*1 = 4 So 4 represents the value of the function at x = 0. Now, thinking about what the function represents (the bouncing ball) what does 4 represent?

Vocaloid:

x = the number of bounces So if x = 0, that means 0 bounces, right? So if 4 is the value of the function at 0 bounces, 4 would be the height at 0 bounces, or the initial height